OpenStudy (anonymous):

a gummy store makes a 13-pound mixture of gummy candy, jelly beans, and hard candy. The cost of gummy candy is $1.00 per pound, jelly beans cost $2.00 per pound, and hard candy costs $2.00 per pound. The mixture calls for three times as many gummy candy pieces as jelly beans. The total cost of the mixture is $20.00. How much of each ingredient did the store use?

hero (hero):

Pounds: gummy + jelly + hard = 13 Dollars: 1gummy + 2jelly + 2hard = 20 Amount: 3jelly = 1gummy

hero (hero):

So 3jelly + jelly + hard = 13 3jelly + 2jelly + 2hard = 20 So 4jelly + 1hard = 13 5jelly + 2hard = 20 So 8jelly + 2hard = 26 5jelly + 2hard = 20 So 3jelly = 6 jelly = 6/3 jelly = 2

OpenStudy (anonymous):

I am confused.

hero (hero):

since 3jelly = 1gummy then 3(2) = gummy `6 = gummy` Now 2 + 6 + hard = 13 8 + hard = 13 hard = 13 - 8 `hard = 5` `jelly = 2`

OpenStudy (anonymous):

Oh I see now. I always have trouble with word problems.

hero (hero):

What are you confused about? The gummy, the jelly or the hard?

OpenStudy (anonymous):

I am not confused now, When you typed out the problem I understood. Word problems just confuse me.

hero (hero):

Really? What about this problem do you understand now? The final answer? Or the steps in between?

OpenStudy (anonymous):

I understand how you got the answer you got.

hero (hero):

Okay, so just to make sure you understand it, would you mind giving a summary of the steps and what was done to produce the final result?

OpenStudy (anonymous):

yes, give me a minute b/c Im about to finish up a diff. prob and then I will write out the steps.

OpenStudy (anonymous):

1. you wrote out eqs. to use 2. Since there are twice as many gummys and jellys you used that and one of the eqs. you wrote to find amount of gummies and jellies 3. you plugged in amount of gummies(6) and amount of jellies (2) to find amount of hard. 4. after finding all amounts you got the answer

hero (hero):

Hmmm, maybe you did understand some of it.

OpenStudy (anonymous):

I did haha I understood a lil on how to do the problem, just needed help fully understanding it. Thank you for helping.

hero (hero):

Did you understand how I set up the initial equations?

OpenStudy (anonymous):

Yes, you got: gummy+jelly+hard=13 b/c the mixture is 13 pounds. You got: 1 gummy + 2 jelly + 2 hard = 20 b/c 1 pound of gummy is 1.00, 1 pound of jelly is 2.00, and 1 pound of hard is 2.00. 20 is the total amount of the mixture. But I am not sure how you got 3 jelly = 1 gummy

OpenStudy (anonymous):

I am stuck on another problem. I have an idea on what to do but I ran into a problem. If I make a post can you help me?